# FairShare: Auditable Geographic Fairness for Multi-Operator LEO Spectrum Sharing

### v1

@@ Abstract
Dynamic spectrum sharing (DSS) among multi-operator low Earth orbit (LEO) mega-constellations is essential for coexistence, yet prevailing policies focus almost exclusively on interference mitigation, leaving geographic equity largely unaddressed. This work investigates whether conventional DSS approaches inadvertently exacerbate the rural digital divide. Through large-scale, 3GPP-compliant non-terrestrial network (NTN) simulations with geographically distributed users, we systematically evaluate standard allocation policies. The results uncover a stark and persistent structural bias: SNR-priority scheduling induces a $1.65\times$ urban-rural access disparity, privileging users with favorable satellite geometry. Counter-intuitively, increasing system bandwidth amplifies rather than alleviates this gap, with disparity rising from $1.0\times$ to $1.65\times$ as resources expand. To remedy this, we propose FairShare, a lightweight, quota-based framework that enforces geographic fairness. FairShare not only reverses the bias, achieving an affirmative disparity ratio of $\Delta_{\text{geo}}=0.72\times$, but also reduces scheduler runtime by 3.3%. This demonstrates that algorithmic fairness can be achieved without trading off efficiency or complexity. Our work provides regulators with both a diagnostic metric for auditing fairness and a practical, enforceable mechanism for equitable spectrum governance in next-generation satellite networks.
This work has been submitted to the IEEE Conference for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ I Introduction
Figure 1: Geographic disparity (a) and the FairShare framework (b). The Physical Problem: Rural users face a structural disadvantage through two compounding mechanisms: beam gain roll-off (as beams are centered on population-dense urban areas) and elevation-dependent path loss. The figure illustrates the latter, where rural users experience lower elevation angles ($\theta_{r}<\theta_{u}$) and longer slant ranges, contributing to the overall SNR penalty (a). The FairShare Solution: The proposed framework replaces purely competition-based allocation with geographic partitioning, enforcing specific bandwidth quotas (b).
@@
This paper addresses this gap through FairShare, which serves both as a benchmarking framework and a geographic-aware allocation policy. Using 3GPP TR 38.811–compliant channel models and simulating large-scale distributed user populations, we systematically evaluate conventional DSS strategies. The results reveal that SNR-priority scheduling induces an urban–rural access disparity of $1.65\times$, while demand-proportional allocation yields a $1.40\times$ gap. Critically, this disparity is shown to be policy-inherent rather than scarcity-driven; increasing bandwidth from 50 to 300 MHz amplifies the priority-based disparity from $1.0\times$ to $1.65\times$.
To overcome this limitation, FairShare introduces explicit, tunable geographic quotas into the allocation process. This simple intervention enforced through a centralized coordination framework representative of emerging regulatory models (see Sec. III) not only eliminates urban bias but achieves affirmative fairness ($\Delta_{\text{geo}}=0.72\times$) while retaining high spectral efficiency. Moreover, the quota-based design reduces computational runtime by 3.3% compared to baseline schedulers.
This work makes four key contributions: (1) an open benchmarking framework for evaluating geographic fairness in the context of multi-operator LEO spectrum sharing under 3GPP channel models; (2) the first systematic evidence that conventional DSS policies induce extreme urban–rural disparities originating from allocation design, not spectrum scarcity; (3) FairShare, a lightweight geographic-aware policy that guarantees equitable access while maintaining high spectral efficiency; and (4) concrete, evidence-based guidance for regulators on incorporating explicit geographic fairness metrics into future NTN spectrum-sharing standards.
@@ II Related Work
Dynamic spectrum sharing in non-terrestrial networks has attracted growing research interest as LEO mega-constellations expand. This section reviews relevant literature across three categories: DSS mechanisms, fairness frameworks, and standardization efforts.
II-A DSS Mechanisms for LEO Satellite Systems
Ahmed et al. [5] provide a comprehensive survey of DSS mechanisms between LEO and terrestrial systems, emphasizing interference management techniques including exclusion zones, power control, and beam coordination. Giordani et al. [11] analyze Ka/Ku-band coexistence challenges in multi-operator settings, identifying regulatory gaps in ITU and FCC frameworks regarding inter-constellation coordination.
Recent artificial intelligence (AI)-driven approaches have emerged for adaptive spectrum allocation. Vazquez et al. [19] propose hierarchical deep reinforcement learning for dynamic spectrum access, achieving improved throughput over rule-based methods. Federated learning approaches have also been explored to enable distributed spectrum decisions without centralized coordination. However, these methods optimize aggregate throughput metrics and embed fairness only implicitly through reward shaping, lacking explicit fairness guarantees or geographic considerations. The opacity of learned policies further presents challenges for regulatory oversight and auditability.
II-B Fairness in Wireless Resource Allocation
Classical fairness metrics Jain’s index [15], proportional fairness, and $\alpha$-fairness [18] were designed for quasi-static terrestrial channels where user conditions are approximately stationary. Ahmad [4] demonstrates that these metrics become unreliable under LEO dynamics: the rapid variation of elevation-dependent SINR causes traditional fairness indices to fluctuate significantly within short time windows, obscuring systematic geographic biases.
Max-min fairness and weighted proportional fairness have been extensively studied in terrestrial cellular networks [8]. These approaches balance efficiency and equity but assume homogeneous user population without geographic structure. When applied to LEO systems where channel quality correlates strongly with location, they fail to detect or prevent geographic discrimination.
II-C Standardization and Regulatory Gaps
@@ III-A Network Topology
III-B Channel Model
@@
$|h_{u}(t)|^{2}=G_{s,b}\cdot G_{u}\cdot L^{-1}(\theta_{u},d_{u})\cdot\xi_{u}(t)\cdot A_{u}(t),$ (1)
where $G_{s,b}$ is the satellite antenna gain, $G_{u}$ is the user terminal gain, $L(\cdot)$ is the elevation-dependent path loss, $\xi_{u}(t)$ is the log-normal shadow fading, and $A_{u}(t)$ is the atmospheric loss [17, Sec. 5.5].
@@
The signal-to-noise-plus-interference ration (SINR) is obtained as
@@
The achievable rate is
@@
III-C Fairness Metrics
@@
where the expectation is taken over channel realizations. For the assumed full-buffer traffic model, unallocated users experience a transmission outage ($R_{u}(t)=0$). Therefore, $\rho_{\ell}$ serves as a direct proxy for service availability, and $(1-\rho_{\ell})$ represents the user outage probability.
@@
to benchmark against traditional notions of fairness We employ this metric in Sec. VII to demonstrate that high aggregate fairness ($J\approx 1$) can paradoxically coexist with severe geographic discrimination.
III-D Problem Formulation
@@
Objective (9) maximizes sum-rate (spectral efficiency) with constraint (9a) to enforce total bandwidth, constraint (9b) to guarantee minimum geographic quotas, and constraint (9c) to ensure minimum allocation rates per region.
@@ IV-A FairShare: Geographic-Aware Allocation
This work is grounded in a guaranteed-minimum, needs-based principle of distributive justice for essential infrastructure. We contend that in broadband access, an immutable geographic disadvantage should not systematically bar users from service. Consequently, we explicitly reject the merit-based principle (exemplified by SNR-priority scheduling) and the strict utilitarian principle (pure throughput maximization) as primary objectives, as they inherently amplify this structural bias. FairShare operationalizes this ethical stance by enforcing minimum geographic quotas, ensuring underserved populations receive a fair opportunity for access.
FairShare defines fairness at the user level, independent of operator identity. Each user’s geographic category determines fairness treatment regardless of which operator serves them, ensuring a rural user receives equitable access whether served by Starlink, OneWeb, or Kuiper.
@@
1:Users $\mathcal{U}$, bandwidth $W$, quotas $\{q_{\ell}\}$, SNR $\{\gamma_{u}\}$
@@
3:for each category $\ell\in\{\text{urban},\text{suburban},\text{rural}\}$ do
4:  $W_{\ell}\leftarrow q_{\ell}\cdot W$ $\triangleright$ Compute quota
5:  $\mathcal{U}_{\ell}\leftarrow\{u:\ell_{u}=\ell\}$ $\triangleright$ Users in region
6:  Sort $\mathcal{U}_{\ell}$ by $\gamma_{u}$ descending
7:  $n_{\ell}\leftarrow\max(1,\lfloor\rho_{\min}\cdot|\mathcal{U}_{\ell}|\rfloor)$ $\triangleright$ Min users
8:  for $u$ in top $n_{\ell}$ of $\mathcal{U}_{\ell}$ do
9:    $b_{u}\leftarrow W_{\ell}/n_{\ell}$
10:  end for
@@ IV-B Optimality Analysis
Setting $q_{\ell}$ proportional to population share yields $\Delta_{\text{geo}}\approx 1.0\times$. The default configuration ($q_{\text{rural}}=35\%$ vs. 30% population share) intentionally over-allocates to rural users, achieving affirmative fairness ($\Delta_{\text{geo}}=0.72\times$).
@@ V-A Simulation Framework and Network Topology
The proposed FairShare policy is evaluated using a high-fidelity, system-level simulation framework built in TensorFlow following 3GPP TR 38.811 channel models, with graphics processing unit (GPU) acceleration. [^3] https://github.com/CLIS-WPI/FairShare We employ a quasi-static Monte Carlo snapshot approach to isolate the impact of structural geometric unfairness. In each simulation instance, satellite positions are fixed to a representative geometry derived from a Starlink-like shell ($S=100$ satellites at 550 km altitude and $53^{\circ}$ inclination) [9]. The propagation channels are generated based on 3GPP TR 38.811 models using spherical earth geometry for precise elevation and slant range calculations.
The network scenario models a multi-operator environment where $O=3$ operators share a common spectrum pool of bandwidth $W=300$ MHz at a carrier frequency of $f_{c}=20$ GHz (Ka-band). User association follows a Best-SNR policy, where each user terminal connects to the visible satellite offering the highest instantaneous channel capacity.
@@ V-B Baseline Policies
V-C Geographic User Distribution
The user distribution is modeled using a concentric statistical framework approximating a typical metropolitan structure with parameters inspired by the New York City urban-rural gradient.
Figure 2: Geographic distribution of 1,000 simulated users. Urban users (50%) cluster near the metropolitan center, while rural users (30%) occupy the outer ring with systematically lower elevation angles.
As illustrated in Fig. 2, rather than relying on static census data, a distance-based classification method is employed to generate a synthetic population gradient. This method operates as follows: (1) Urban users (50% of the total population) are clustered via a Gaussian distribution ($\sigma\approx 0.05^{\circ}\approx 5.5$ km) centered at the metropolitan core; (2) Suburban users (20%) are uniformly distributed within an annular ring at 22–55 km from the center; and (3) Rural users (30%) are distributed in an outer ring extending from 55–165 km. These distances are consistent with typical metropolitan population density gradients and ensure a continuous variation of satellite elevation angles and path loss conditions.
V-D Channel and Traffic Models
Channel propagation is modeled according to 3GPP TR 38.811 specifications [2]. The link budget assumes a satellite EIRP of 45 dBW, a user terminal gain of 30 dBi, and a noise figure of 2 dB, representative of Ka-band LEO systems [3, 16]. Shadow fading is location-dependent, with standard deviations set to $\sigma_{\text{SF}}=8$ dB for urban and $\sigma_{\text{SF}}=4$ dB for rural environments, reflecting the higher obstacle density in city centers. Regarding traffic dynamics, we assume a full-buffer downlink traffic model where demand is constant across all time slots ($T=100$). This assumption allows evaluating the system under worst-case contention scenarios where spectrum scarcity is most pronounced.
Statistical Methodology: To ensure statistical robustness, each policy is evaluated over $n=50$ independent Monte Carlo runs. The low empirical variance confirms this sample size is sufficient for stable estimates. Results are reported as mean $\pm$ standard deviation. The expectation in (5) is estimated via Monte Carlo averaging.

### v2

@@ Abstract
Dynamic spectrum sharing (DSS) among multi-operator low Earth orbit (LEO) mega-constellations is essential for coexistence, yet prevailing policies focus almost exclusively on interference mitigation, leaving geographic equity largely unaddressed. This work investigates whether conventional DSS approaches inadvertently exacerbate the rural digital divide. Incorporating Keplerian orbital dynamics, inter-beam co-channel interference, and three real-world constellation geometries (Starlink, OneWeb, Kuiper), we conduct large-scale, 3GPP-compliant non-terrestrial network (NTN) simulations across 20 orbital snapshots spanning 10 minutes of satellite motion. The results uncover a stark and persistent structural bias: SNR-priority scheduling induces a $1.84\times$ mean urban–rural access disparity, with temporal fluctuations reaching $3.9\times$ during favorable interference conditions. Counter-intuitively, increasing system bandwidth amplifies rather than alleviates this gap. To remedy this, we propose FairShare, a lightweight, quota-based framework that enforces geographic fairness. FairShare not only reverses the bias, achieving an affirmative disparity ratio of $\Delta_{\text{geo}}=0.68\times$ with zero variance across all orbital snapshots and interference conditions, but also reduces scheduler runtime by 3.3%. This demonstrates that algorithmic fairness can be achieved without trading off efficiency or complexity, and that it remains invariant to physical-layer dynamics. Our work provides regulators with both a diagnostic metric for auditing fairness and a practical, enforceable mechanism for equitable spectrum governance in next-generation satellite networks.
@@ I Introduction
Figure 1: Geographic disparity (a) and the FairShare framework (b). The Physical Problem: Rural users face a structural disadvantage through two compounding mechanisms: beam gain roll-off (as beams are centered on population-dense urban areas) and elevation-dependent path loss. The figure illustrates the latter, where rural users experience lower elevation angles ($\theta_{r}<\theta_{u}$) and longer slant ranges, contributing to the overall SNR penalty (a). The FairShare Solution: The proposed framework replaces purely competition-based allocation with geographic partitioning, enforcing specific bandwidth quotas (b). Note: the SNR gap shown is illustrative; measured values are 5–10 dB.
@@
This paper addresses this gap through FairShare, which serves both as a benchmarking framework and a geographic-aware allocation policy. Using 3GPP TR 38.811 compliant channel models and simulating large-scale distributed user populations, we systematically evaluate conventional DSS strategies. The results reveal that SNR-priority scheduling induces a mean urban rural access disparity of $1.84\times$, with temporal fluctuations reaching $3.9\times$ during favorable interference conditions, while demand-proportional allocation yields a $1.31\times$ gap. Critically, this disparity is shown to be policy-inherent rather than scarcity-driven and persists across all three constellation geometries evaluated.
To overcome this limitation, FairShare introduces explicit, tunable geographic quotas into the allocation process. This simple intervention enforced through a centralized coordination framework representative of emerging regulatory models (see Sec. III) not only eliminates urban bias but achieves affirmative fairness ($\Delta_{\text{geo}}=0.68\times$) with zero temporal variance, while retaining high spectral efficiency. Moreover, the quota-based design reduces computational runtime by 3.3% compared to baseline schedulers.
This work makes five key contributions: (1) an open benchmarking framework for evaluating geographic fairness in multi-operator LEO spectrum sharing under 3GPP channel models with Keplerian orbital propagation and inter-beam co-channel interference; (2) the first systematic evidence that conventional DSS policies induce extreme urban-rural disparities originating from allocation design, not spectrum scarcity, validated across 20 orbital snapshots spanning 10 minutes of satellite motion; (3) FairShare, a lightweight geographic-aware policy that guarantees equitable access while maintaining high spectral efficiency; (4) cross-constellation validation demonstrating that the structural bias persists and FairShare remains effective across Starlink, OneWeb, and Kuiper geometries; and (5) concrete, evidence-based guidance for regulators on incorporating explicit geographic fairness metrics into future NTN spectrum-sharing standards.
@@ II Related Work
II-A DSS and Fairness in LEO Systems
Ahmed et al. [5] survey DSS mechanisms between LEO and terrestrial systems, emphasizing interference management, while Giordani et al. [11] analyze Ka/Ku-band coexistence challenges in multi-operator settings, identifying regulatory gaps regarding inter-constellation coordination. Recent AI-driven approaches such as hierarchical deep reinforcement learning [20] achieve improved throughput but optimize aggregate metrics, lacking explicit fairness guarantees or geographic considerations; the opacity of learned policies further presents challenges for regulatory auditability.
Classical fairness metrics Jain’s index [16], proportional fairness, $\alpha$-fairness [19] were designed for quasi-static terrestrial channels. Ahmad [4] demonstrates that these become unreliable under LEO dynamics, where elevation-dependent SINR fluctuations obscure systematic geographic biases. Max-min and weighted proportional fairness [8] assume homogeneous user populations without geographic structure and fail to detect location-correlated discrimination.
II-B Standardization and Regulatory Gaps
@@ III-A Network Topology
III-B Orbital Dynamics and Constellation Model
To evaluate fairness under realistic satellite motion, we employ a Keplerian orbital propagator. Each satellite’s position in Earth-Centered Inertial (ECI) coordinates is computed from classical two-body orbital elements and transformed to Earth-Centered Earth-Fixed (ECEF) coordinates accounting for Earth’s rotation:
$\mathbf{r}_{\text{ECEF}}(t)=\mathbf{R}_{z}(\omega_{E}t)\,\mathbf{r}_{\text{ECI}}(t),$ (1)
where $\omega_{E}=7.2921\times 10^{-5}$ rad/s is Earth’s rotation rate and $\mathbf{R}_{z}(\cdot)$ is the rotation matrix about the polar axis. The simulation runs $N_{\text{snap}}=20$ orbital snapshots at $\Delta t=30$ s intervals, spanning 10 minutes of satellite motion. At each snapshot, satellite positions are re-propagated and user satellite geometry (elevation angles, slant ranges) is recomputed. Only satellites exceeding a minimum elevation angle $\theta_{\min}=10^{\circ}$ (per 3GPP TR 38.811 [2]) are considered visible.
Table I defines the three Walker-Delta constellation configurations evaluated. These span the range of current LEO filings in altitude (550–1,200 km), inclination (51.9–87.9), and constellation size (648–1,584 satellites).
TABLE I: Constellation configurations evaluated.
Constellation Planes Sats/Pl. Total Alt. Inc.
Starlink Shell 1 72 22 1,584 550 km 53.0
OneWeb Phase 1 18 36 648 1,200 km 87.9
Kuiper Shell 1 34 34 1,156 630 km 51.9
III-C Channel Model
@@
$|h_{u}(t)|^{2}=G_{s,b}(\psi_{u})\cdot G_{u}\cdot L^{-1}(\theta_{u},d_{u})\cdot\xi_{u}(t)\cdot A_{u}(t),$ (2)
where $G_{s,b}(\psi_{u})$ is the satellite beam gain as a function of the user’s off-axis angle $\psi_{u}$ from the serving beam center (Sec. III-D), $G_{u}$ is the user terminal gain, $L(\cdot)$ is the elevation-dependent path loss, $\xi_{u}(t)$ is the log-normal shadow fading, and $A_{u}(t)$ is the atmospheric loss [18, Sec. 5.5].
@@
The signal-to-interference-plus-noise ratio (SINR) is obtained as
@@
where $p_{s,b,u}(t)$ denotes the transmit power allocated from satellite $s$ and beam $b$ to user $u$ at time $t$, $I_{u}^{(s^{\prime},b^{\prime})}(t)$ is the interference power received at user $u$ from beam $b^{\prime}$ of satellite $s^{\prime}$, $b_{u}(t)$ is the bandwidth allocated to user $u$ (Hz), and $N_{0}$ is the noise power spectral density so that $N_{0}b_{u}(t)$ is the corresponding noise power. The achievable rate is
@@
III-D Beam Model and Interference
Each satellite generates a hexagonal 7-beam spot-beam layout centered on its sub-satellite point. The antenna gain for user $u$ at angular offset $\psi_{u}$ from beam center follows the ITU-R S.1528 [14] parabolic approximation:
$G_{s,b}(\psi_{u})=G_{\max}-12\!\left(\frac{\psi_{u}}{\psi_{3\text{dB}}}\right)^{\!2}\quad[\text{dBi}],$ (6)
clamped at $G_{\max}-25$ dB (first sidelobe floor), where $G_{\max}=30$ dBi is the peak beam gain and $\psi_{3\text{dB}}=1.5^{\circ}$ is the half-power beamwidth for Ka-band spot beams. A 4-color frequency reuse pattern is applied across the 7 beams, assigning each beam one of four sub-bands such that adjacent beams operate on distinct frequencies.
Co-channel interference arises from beams sharing the same sub-band. The interferer set for user $u$ served by beam $b$ on frequency $f_{b}$ is
$\mathcal{I}_{u}=\{(s^{\prime},b^{\prime}):f_{b^{\prime}}=f_{b},\;(s^{\prime},b^{\prime})\neq(s,b)\}.$ (7)
The 4-color reuse significantly reduces interference by ensuring that only $\sim$25% of all beams are co-channel with the serving beam. The resulting SINR is 15–25 dB lower than the interference-free SNR, placing users in a realistic operating regime for Ka-band LEO systems.
III-E Fairness Metrics
@@
where the expectation is taken over channel realizations, and $\mathcal{U}_{\ell}=\{u:\ell_{u}=\ell\}$ denotes the set of users in geographic category $\ell$. For the assumed full-buffer traffic model, unallocated users experience a transmission outage ($R_{u}(t)=0$). Therefore, $\rho_{\ell}$ serves as a direct proxy for service availability, and $(1-\rho_{\ell})$ represents the user outage probability.
@@
to benchmark against traditional notions of fairness. We employ this metric in Sec. VII to demonstrate that high aggregate fairness ($J\approx 1$) can paradoxically coexist with severe geographic discrimination.
III-F Problem Formulation
@@
For notational simplicity, we drop the explicit time index and write $b_{u}$ and $\gamma_{u}$ as representative per-slot bandwidth allocation and SINR, respectively. Objective (12) maximizes sum-rate (spectral efficiency) with constraint (12a) to enforce total bandwidth, constraint (12b) to guarantee minimum geographic quotas, and constraint (12c) to ensure minimum allocation rates per region.
@@ IV-A FairShare: Geographic-Aware Allocation
FairShare is grounded in a guaranteed-minimum principle: an immutable geographic disadvantage should not systematically bar users from service. It operationalizes this by enforcing minimum geographic quotas, rejecting pure merit-based (SNR-priority) and utilitarian (throughput-maximizing) objectives that amplify structural bias. Fairness is defined at the user level, independent of operator identity, ensuring equitable access regardless of which constellation serves a given user.
@@
1:Users $\mathcal{U}$, bandwidth $W$, quotas $\{q_{\ell}\}$, SINR $\{\gamma_{u}\}$, min per-user bandwidth $b_{\min}$
@@
[…]
4:for each category $\ell\in\{\text{urban},\text{suburban},\text{rural}\}$ do
5:  $W_{\ell}\leftarrow q_{\ell}\cdot W$ $\triangleright$ Region bandwidth quota
6:  $\mathcal{U}_{\ell}\leftarrow\{u:\ell_{u}=\ell\}$ $\triangleright$ Users in region
7:  $n_{\ell}\leftarrow\min\!\bigl(\max(1,\lfloor N_{\text{alloc}}\cdot q_{\ell}\rfloor),\;|\mathcal{U}_{\ell}|\bigr)$ $\triangleright$ User slots for region
8:  Sort $\mathcal{U}_{\ell}$ by $\gamma_{u}$ descending
9:  for $u$ in top $n_{\ell}$ of $\mathcal{U}_{\ell}$ do
10:    $b_{u}\leftarrow W_{\ell}/n_{\ell}$
@@ IV-B Optimality Analysis
Setting $q_{\ell}$ proportional to population share yields $\Delta_{\text{geo}}\approx 1.0\times$. The default configuration ($q_{\text{rural}}=35\%$ vs. 30% population share) intentionally over-allocates to rural users, achieving affirmative fairness ($\Delta_{\text{geo}}=0.68\times$).
@@ V-A Simulation Framework and Network Topology
The proposed FairShare policy is evaluated using a high-fidelity, system-level simulation framework built in TensorFlow following 3GPP TR 38.811 channel models, with graphics processing unit (GPU) acceleration. [^3] https://github.com/CLIS-WPI/FairShare We employ a multi-snapshot Monte Carlo approach incorporating Keplerian orbital propagation (Sec. III-B). The simulation runs $N_{\text{snap}}=20$ snapshots at 30-second intervals, spanning 10 minutes of satellite motion. At each snapshot, satellite positions are re-propagated and the full channel model including inter-beam co-channel interference (Sec. III-D) is recomputed. Three Walker-Delta constellation configurations are evaluated (Table I): Starlink Shell 1 (1,584 satellites at 550 km), OneWeb Phase 1 (648 satellites at 1,200 km), and Kuiper Shell 1 (1,156 satellites at 630 km) [9].
The coverage area is centered at the New York City metropolitan area ($40.7^{\circ}$N, $74.0^{\circ}$W). The network scenario models a multi-operator environment where $O=3$ operators share a common spectrum pool of bandwidth $W=300$ MHz at a carrier frequency of $f_{c}=20$ GHz (Ka-band). User association follows a Best-SINR policy, where each user terminal connects to the visible satellite offering the highest instantaneous channel quality.
@@ V-B Baseline Policies
V-C Geographic User Distribution and Channel Model
Figure 2: Geographic distribution of 1,000 simulated users centered at NYC ($40.7^{\circ}$N, $74.0^{\circ}$W). Urban (50%), suburban (20%), rural (30%).
Users are classified by distance from the metropolitan center (Fig. 2): Urban (50%, Gaussian $\sigma\approx 5.5$ km), Suburban (20%, 22–55 km annulus), Rural (30%, 55–165 km ring). Channel propagation follows 3GPP TR 38.811 [2] with EIRP == 45 dBW, user terminal gain == 30 dBi, noise figure == 2 dB [3, 17], and location-dependent shadow fading ($\sigma_{\text{SF}}=8$ dB urban, 4 dB rural). A full-buffer downlink traffic model ($T=100$ slots per snapshot) provides a worst-case contention baseline.
Statistical Methodology: Each policy is evaluated over $20\times 50=1{,}000$ independent samples per constellation. Standard deviations reflect temporal variability due to orbital dynamics rather than statistical uncertainty.